# MORDA: A Synthetic Dataset to Facilitate Adaptation of Object Detectors to Unseen Real-target Domain While Preserving Performance on Real-source Domain

### v1

@@ Abstract
Deep neural network (DNN) based perception models are indispensable in the development of autonomous vehicles (AVs). However, their reliance on large-scale, high-quality data is broadly recognized as a burdensome necessity due to the substantial cost of data acquisition and labeling. Further, the issue is not a one-time concern, as AVs might need a new dataset if they are to be deployed to another region (real-target domain) that the in-hand dataset within the real-source domain cannot incorporate. To mitigate this burden, we propose leveraging synthetic environments as an auxiliary domain where the characteristics of real domains are reproduced. This approach could enable indirect experience about the real-target domain in a time- and cost-effective manner. As a practical demonstration of our methodology, nuScenes and South Korea are employed to represent real-source and real-target domains, respectively. That means we construct digital twins for several regions of South Korea, and the data-acquisition framework of nuScenes is reproduced. Blending the aforementioned components within a simulator allows us to obtain a synthetic-fusion domain in which we forge our novel driving dataset, MORDA: Mixture Of Real-domain characteristics for synthetic-data-assisted Domain Adaptation. To verify the value of synthetic features that MORDA provides in learning about driving environments of South Korea, 2D/3D detectors are trained solely on a combination of nuScenes and MORDA. Afterward, their performance is evaluated on the unforeseen real-world dataset (AI-Hub111This research (paper) used datasets from High-precision data collection vehicle daytime city road data. All data information can be accessed through AI-Hub (http://www.aihub.or.kr ).) collected in South Korea. Our experiments present that MORDA can significantly improve mean Average Precision (mAP) on AI-Hub dataset while that on nuScenes is retained or slightly enhanced. Details on MORDA will be available at https://morda-e8d07e.gitlab.io.
@@ I INTRODUCTION
To alleviate the labeling cost, research on unsupervised domain adaptation (UDA) has been actively conducted. UDA methods assume that only the is labeled, and they exploit the knowledge from the annotated for better adaptation (i.e., performance) of DNNs to the unlabeled without additional labeling costs [11, 12]. Nevertheless, dispatching the sensor-equipped vehicle to for raw sensor data collection is inevitable. Furthermore, the engineering cost for synchronizing raw data from sensors operating at different frequencies is non-negligible.
[b] Sensor setup BBox annotation Camera* Resolution LiDAR Channel Sensor suite reference DOF† Road obstacle‡ Frequency(Hz) VIPER [13] 1 19201080 ✗ ✗ - 9 ✗ ~15 VKITTI2 [14] 5 1242375 ✗ ✗ KITTI [15] 9 ✗ - SYNTHIA-AL [16] 1 640480 ✗ ✗ - 7 ✗ 25 SHIFT [17] 5 1200800 1 128 - 9 ✗ 10 MORDA (Ours) 6 1600900 1 32 nuScenes [18] 9 ✓ 20
@@
Degree of freedom, 7 and 9 indicates the orientation is represented with (yaw) and (roll, pitch, and yaw), respectively.
@@
Given the mentioned constraints in UDA and existing synthetic datasets, the research objective of this paper is to develop a synthetic dataset that enhances the adaptation of DNNs to while their performance on is not degraded. Note that this scenario is challenging as (1) is completely unforeseen, meaning sensor data of the target domain is not used for training DNNs, and (2) the synthetic dataset should improve the generalization ability of DNNs as they need to perform well on both and .
We believe an adequate synthetic dataset for our purpose needs to be equipped with the following attributes; (1) Virtual version of geographical features of driving environments in the target domain (), which enables cost-effective indirect exposure of to DNNs. (2) Consistency with in terms of sensor suite and labeling process to keep the characteristics of created synthetic dataset paired with . In this regard, we propose a methodology to generate a synthetic dataset in a synthetic-fusion domain () where the aforementioned aspects of and are well integrated. As a demonstration, we construct a novel synthetic dataset, MORDA which takes nuScenes for and South Korea for . Next, we evaluate the efficacy of our MORDA dataset in a challenging scenario, as illustrated in Fig. 1. Firstly, 2D/3D object detectors are trained solely on nuScenes () combined with MORDA (). Afterward, they are evaluated on the unforeseen AI-Hub dataset collected in South Korea (). Our extensive experiments present that MORDA allows notable detection-performance gain on while successfully maintaining the performance on simultaneously. In short, the main contributions are:
@@
Using MORDA, we show that utilizing could be a financially-viable method for object detectors to adapt without performance degradation in .
@@ II-A Object Detection
The field of object detection has shown surprising progress with DNNs in recent years, which aims to the accurate prediction of bounding box (BBox) to localize and classify objects within the sensor’s field-of-view. The prosperity of DNN-based object detection applies to not only 2D but also challenging 3D detection tasks across various sensor modalities [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. However, the inherent challenge of learning algorithms remains that performance degradation occurs when DNNs are evaluated on a new domain which is unforeseen during training [21, 17]. Note that addressing such is imperative for AVs whose driving environments or conditions can often change, e.g., deploying the developed AVs to another country ().
@@ II-B Unsupervised Domain Adaptation (UDA)
To tackle the performance drop in new domains, numerous studies have concentrated on adapting DNNs trained on the source domain (w/ labels) to the target domain (w/o labels). Recent works in this field present significant advances even in challenging LiDAR-based 3D object detection. For example, [11, 12, 22] address the adaptation problem of Waymo [23] nuScenes [18], presenting notable improvement. Nevertheless, several drawbacks remain that (1) UDA often requires a sophisticated training strategy, e.g., adversarial training scheme [24, 22] and multiple rounds of self-training [25, 26]. (2) Potential risk of error propagation exists as many UDA methods rely on pseudo labeling [25, 26, 11, 12]. (3) Sensor data from the target domain, although not need to be labeled, are required which could be still expensive in practice.
@@ II-C Synthetic Driving Datasets
Cost-efficient virtual environments have been actively employed to reduce the burden of constructing large-scale driving datasets with rich annotations. For example, Unity222http://unity3d.com/ engine is used to generate Virtual KITTI 2 (VKITTI2) [14], and SYNTHIA-AL [16]. A modern game Grand Theft Auto V is utilized to create GTA5 [19] and VIPER [13]. Recently, open-source driving simulator CARLA [27] has been employed to produce SHIFT [17].
@@ III-A Overview
The purpose of our method is to construct a dataset from a synthetic-fusion domain () where DNNs (specifically, 2D/3D object detectors) can indirectly learn about the unforeseen to enhance adaptability while retaining the performance on at which the model is trained. To accomplish the aim, we extract pivotal characteristics from and . Then, they are reproduced and blended within a simulator, creating our fusion domain. Lastly, our synthetic dataset (MORDA) is generated using the simulator. Fig. 2 illustrates this procedure in detail. As a showcase of the proposed methodology, we choose nuScenes and South Korea to represent and , respectively. Next, the MORAI Simulator (MORAI SIM) [28] is employed to provide a baseline simulation environment where fundamental features like driving agents and GT-label generation are available.
@@ IV-A Dataset Content
MORDA comprises ~37K frames where each frame consists of ego-vehicle pose data, six 1600900 images, and one 32-channel point cloud data with corresponding GT labels. The types of GT include pixel/point-level semantic and instance segmentation, 2D/3D BBox annotation, and pixel-wise depth values. All types of GT labels are generated by individual virtual sensors independently. Therefore, our dataset can be utilized for various vision tasks such as monocular/multi-view camera-based, lidar-based, fusion-based object detection, and segmentation tasks.
IV-B Comaprison to nuScenes Dataset
@@ V-A3 Metrics
We use renowned average precision (AP), mean average precision (mAP), and nuScenes detection score (NDS) [18]. All metrics are represented in percentage (), and the higher the better.
@@ V-B1 Real-target dataset
A real-world driving dataset, AI-Hub is opted to represent given (1) its sensor data are collected from the real-target domain, South Korea where digital twin maps in MORDA () is originated from. (2) Sensor suite and GT types are similar to nuScenes and MORDA, i.e., each frame contains five 19201200 multi-view images, one point cloud from 128-channel spinning LiDAR, and RTK GNSS data with 3D-BBox annotation frequency of 10 Hz.
@@ V-C1 Pre-processing
We use images from the front camera that all mentioned datasets above have in common. Next, we sample the datasets to match their frequencies as much as possible. That means MORDA, SYNTHIA-AL, and VKITTI2444The frequency of VKITTI2 is assumed to be the same as KITTI, as we could not find relevant information. are downsampled to 2Hz. Although images of SHIFT dataset (discrete-shift) are already of 1 Hz, we downsample it by a factor of 5 due to its exceptional size (150K). Note that SHIFT is still the largest dataset after this downsampling, as shown in Tab. IV. 2D BBoxes of respective datasets are merged to nuScenes by class labels.
@@ VI CONCLUSIONS
For safety-critical AVs, DNNs need to overcome the performance degradation that occurs in unexperienced while maintaining performance on . This paper proposed leveraging where key characteristics of and are fused. To showcase the efficacy, we generated a novel dataset, MORDA. Our comprehensive experiments demonstrate that MORDA could help DNNs adapt to , mitigating their dependency on real-world data without losing performance on . Nevertheless, the potential of MORDA is not fully explored as only the training strategy of simple dataset concatenation was applied. Thus, future work could be investigating sophisticated training methods to further enhance the performance on .

### v2

@@ Abstract
Deep neural network (DNN) based perception models are indispensable in the development of autonomous vehicles (AVs). However, their reliance on large-scale, high-quality data is broadly recognized as a burdensome necessity due to the substantial cost of data acquisition and labeling. Further, the issue is not a one-time concern as AVs might need a new dataset if they are to be deployed to another region (real-target domain) that the in-hand dataset within the real-source domain cannot incorporate. To mitigate this burden, we propose leveraging synthetic environments as an auxiliary domain where the characteristics of real domains are reproduced. This approach could enable indirect experience about the real-target domain in a time- and cost-effective manner. As a practical demonstration of our methodology, nuScenes and South Korea are employed to represent real-source and real-target domains, respectively. That means we construct digital twins for several regions of South Korea, and the data-acquisition framework of nuScenes is reproduced. Blending the aforementioned components within a simulator allows us to obtain a synthetic-fusion domain in which we forge our novel driving dataset, MORDA: Mixture Of Real-domain characteristics for synthetic-data-assisted Domain Adaptation. To verify the value of synthetic features that MORDA provides in learning about driving environments of South Korea, 2D/3D detectors are trained solely on a combination of nuScenes and MORDA. Afterward, their performance is evaluated on the unforeseen real-world dataset (AI-Hub111This research (paper) used datasets from High-precision data collection vehicle daytime city road data. All data information can be accessed through AI-Hub (http://www.aihub.or.kr ).) collected in South Korea. Our experiments present that MORDA can significantly improve mean Average Precision (mAP) on AI-Hub dataset while that on nuScenes is retained or slightly enhanced. Details on MORDA can be accessed at https://morda-e8d07e.gitlab.io.
@@ I INTRODUCTION
To alleviate the labeling cost, research on unsupervised domain adaptation (UDA) has been actively conducted. UDA methods assume that only the is labeled, and they exploit the knowledge from the annotated for better adaptation (i.e., performance) of DNNs to the unlabeled without additional labeling costs [11, 12]. Nevertheless, dispatching the sensor-equipped vehicle to for data collection is inevitable even in this scenario. Further, the engineering cost to synchronize raw data from sensors operating often at different frequencies is non-negligible.
[b] Sensor setup BBox annotation Camera* Resolution LiDAR Channel Sensor suite reference DOF† Road obstacle‡ Frequency(Hz) VIPER [13] 1 19201080 ✗ ✗ - 9 ✗ ~15 VKITTI2 [14] 5 1242375 ✗ ✗ KITTI [15] 9 ✗ - SYNTHIA-AL [16] 1 640480 ✗ ✗ - 7 ✗ 25 SHIFT [17] 5 1200800 1 128 - 9 ✗ 10 OPV2V[18] 4 800600 1 64 - 9 ✗ 10 MORDA (Ours) 6 1600900 1 32 nuScenes [19] 9 ✓ 20
@@
Degree of freedom, 9 means 3D-BBox labels are represented with position (x, y, z), size (length, width, height), and orientation (roll, pitch, yaw). 7 denotes the orientation is represented only with yaw angle.
@@
Given the aforementioned constraints in UDA and existing synthetic datasets, the research objective of this paper is to develop a synthetic dataset that enhances the adaptation of DNNs to while their performance on is not degraded. Note that this scenario is challenging as (1) is completely unforeseen, meaning sensor data of the target domain is not used for training DNNs, and (2) the synthetic dataset should improve the generalization ability of DNNs as they need to perform well on both and .
We believe an adequate synthetic dataset for our purpose needs to be equipped with the following attributes; (1) Virtual version of geographical features of driving environments in the target domain (), which enables cost-effective indirect exposure of to DNNs. (2) Consistency with in terms of sensor suite and labeling process to keep the characteristics of created synthetic dataset paired with . In this regard, we propose a methodology to generate a synthetic dataset in a synthetic-fusion domain () where the mentioned aspects of and are well integrated. As a demonstration, we construct a novel synthetic dataset, MORDA which takes nuScenes for and South Korea for . Next, we evaluate the efficacy of our MORDA dataset in a challenging scenario, as illustrated in Fig. 1. Firstly, 2D/3D object detectors are trained solely on nuScenes () combined with MORDA (). Afterward, they are evaluated on the unforeseen driving scenes of South Korea () present in AI-Hub dataset. Our extensive experiments demonstrate that MORDA allows noteworthy detection-performance gain on while successfully maintaining the performance on simultaneously. In short, the main contributions are:
@@
Using MORDA, we show that utilizing could be a financially-viable method for detectors to adapt to without performance degradation on .
@@ II-A Object Detection
The field of object detection has shown surprising progress with DNNs in recent years, which aims to the accurate prediction of bounding box (BBox) to localize and classify objects within the sensor’s field-of-view. The prosperity of DNN-based object detection applies to not only 2D but also challenging 3D detection tasks across various sensor modalities [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. However, the inherent challenge of learning algorithms remains that the performance degradation occurs when DNNs are evaluated on a new domain which is unforeseen during training [22, 17]. Note that addressing such is imperative for AVs whose driving environments or conditions can often change, e.g., deploying the developed AVs to another country ().
@@ II-B Unsupervised Domain Adaptation (UDA)
To tackle the performance drop in new domains, numerous studies have concentrated on adapting DNNs trained on the source domain (w/ labels) to the target domain (w/o labels). Recent works in this field present notable advances even in challenging LiDAR-based 3D object detection. For example, [11, 12, 23] address the adaptation problem of Waymo [24] nuScenes [19], presenting notable improvement.
Despite the striving of UDA, several difficulties remain to be resolved. Firstly, adversarial training schemes are often employed to extract domain-invariant features, expecting the learned representations to adapt seamlessly to the target domain [25, 23]. However, as noted in [26], if there is a severe distribution mismatch between source and target domains, it can lead to insufficient generalization. Secondly, sophisticated self-training schemes [27, 28, 11, 12], which leverage knowledge obtained from the labeled source domain to generate pseudo labels for the target domain, naturally risk potential error propagation. Lastly, sensor data collection at the target regions is still required, which could be expensive in practice even without labeling.
@@ II-C Synthetic Driving Datasets
Cost-efficient virtual environments have been actively employed to reduce the burden of constructing large-scale driving datasets with rich annotations. For example, Unity222http://unity3d.com/ engine is used to generate Virtual KITTI 2 (VKITTI2) [14], and SYNTHIA-AL [16]. A modern game Grand Theft Auto V is utilized to create GTA5 [20] and VIPER [13]. Recently, open-source driving simulator CARLA [29] has been employed to produce SHIFT [17] and OPV2V [18].
@@ III-A Overview
The purpose of our method is to construct a dataset from a synthetic-fusion domain () where DNNs (specifically, 2D/3D object detectors) can indirectly learn about the unforeseen to enhance adaptability while retaining the performance on at which the model is trained. To accomplish the aim, we extract pivotal characteristics from and . Then, they are reproduced and blended within a simulator, creating our fusion domain. Lastly, our synthetic dataset (MORDA) is generated using the simulator. Fig. 2 illustrates this procedure in detail. As a showcase of the proposed methodology, we choose nuScenes and South Korea to represent and , respectively. Next, the MORAI Simulator (MORAI SIM) [30] is employed to provide virtual environment where fundamental features like driving agents and GT-label generation are available.
@@ IV-A Dataset Content
MORDA comprises ~37K frames where each frame consists of ego-vehicle pose data, six 1600900 images, and one 32-channel point cloud data with corresponding GT labels. The types of GT include pixel/point-level semantic and instance segmentation, 2D/3D BBox annotation, and pixel-wise depth values. All types of GT labels are generated independently by individual virtual sensors. Therefore, our dataset can be utilized for various vision tasks such as monocular/multi-view camera-based, lidar-based, fusion-based object detection, and segmentation tasks.
IV-B Comparison to nuScenes Dataset
@@ V-A3 Metrics
We use renowned average precision (AP), mean average precision (mAP), and nuScenes detection score (NDS) [19]. All metrics are represented in percentage (), and higher scores indicate better performance.
@@ V-B1 Real-target dataset
A real-world driving dataset, AI-Hub is opted to represent given (1) its sensor data are collected from the real-target domain of this paper, South Korea where digital twin maps in MORDA () are originated from. (2) Sensor suite and GT types are similar to nuScenes and MORDA, i.e., each frame contains five 19201200 multi-view images, one point cloud from 128-channel spinning LiDAR, and RTK GNSS data with 3D-BBox annotation frequency of 10 Hz.
@@ V-C1 Pre-processing
We use images from the front camera that all mentioned datasets above have in common. Next, we sample the datasets to match their frequencies as much as possible. That means MORDA, SYNTHIA-AL, and VKITTI2444The frequency of VKITTI2 is assumed to be the same as KITTI, as we could not find relevant information. are downsampled to 2Hz. Although images of SHIFT dataset (discrete-shift) are already of 1 Hz, we downsample it by a factor of 5 due to its exceptional size (150K). Note that SHIFT is still the largest dataset after this downsampling, as shown in Tab. IV. 2D BBoxes of respective datasets are merged to nuScenes based on class labels.
@@ VI CONCLUSIONS
For safety-critical AVs, DNNs need to overcome the performance degradation that occurs in unexperienced while maintaining performance on . This paper proposed leveraging where key characteristics of and are fused. To showcase the efficacy, we generated a novel synthetic dataset, MORDA. Our comprehensive experiments demonstrate that MORDA could help DNNs adapt to , mitigating their dependency on real-world data without compromising performance on .
Nevertheless, there is room for further research. Since only a simple concatenation of nuScenes and MORDA was used in this study, further performance enhancement on might be achieved by additionally incorporating modern UDA methods. Another direction for future work could be to identify which characteristics implemented in MORDA contributed to performance stability on by conducting an in-depth comparison with other synthetic datasets.